OpenStudy (hugy0212):

Stanford University's soccer field has an area of 8800 square yards. Its length is 30 yards longer than its width. 1. Write an expression for each dimension (length and width) of the soccer field. 2.Write an equation and solve for the dimensions of the soccer field.

Vocaloid (vocaloid):

let length = L and width = W, with me so far?

OpenStudy (hugy0212):

yes

Vocaloid (vocaloid):

"length is 30 yards longer than its width. " means that L = 30 + W so length = 30 + W and width = W, that takes care of #1

Vocaloid (vocaloid):

for #2, we let length*width = area so W(W+30) = 8800 solve for W

OpenStudy (hugy0212):

Do I distribute the W ?

Vocaloid (vocaloid):

yes

OpenStudy (hugy0212):

W^2+30W=8800

Vocaloid (vocaloid):

good, now subtract 8800 from each side

OpenStudy (hugy0212):

Is it -8800W^2-8770W?

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

remember that you cannot combine numbers and variables

Vocaloid (vocaloid):

W^2+30W-8800=0

OpenStudy (hugy0212):

So it'll be W^2+30W-8800

Vocaloid (vocaloid):

yes, now factor

Vocaloid (vocaloid):

W^2+30W-8800 = ?

OpenStudy (hugy0212):

can and 10 and 880 work?

Vocaloid (vocaloid):

nope, think carefully our two factors must add to 30 and multiply to get 8800

Vocaloid (vocaloid):

W^2+30W-8800 = (W+110)(W-80), know where to go from here?

OpenStudy (hugy0212):

Do i solve for W?

Vocaloid (vocaloid):

yes

OpenStudy (hugy0212):

W=-110 and W=80

Vocaloid (vocaloid):

good, now remember that W is width, and width must be positive so W = 80

Vocaloid (vocaloid):

now find L

OpenStudy (hugy0212):

L=110

Vocaloid (vocaloid):

good! that should b everything

OpenStudy (hugy0212):

Thank you so much!!